UT 20140119 LUCI2 Nighttime Commissioning

Observer: JHeidt, APramskiy, MEdwards
Telescope Operator: SAllenson
LBTO Support: TSargent

Plan for the night

Check for CofR
LS-alignment 2 stars with small rotation angle
Tests scripting binocular mode
Tests in prep of active flexure compensation; short spectra with flex.comp off
Simulating scientific observations
Since number of errors/bugs has decreased considerably now, we will emphasize them

Summary

- GEIRS again stable all night

- SW stable too, except still persistent problem with telescope service
(will be solved tomorrow, TSargent has provided the finite solution)

- Simulating science observations (deep nB-imaging, deep spectroscopy)

- Some tests in binocular mode

- Otherwise smooth night

Preparations

Dave did again calibration for the previous night in the morning
Field alignment checked: dx=-1100 steps, dy=-3800steps John/Doug were working on the collimation issues from last night, seems that they have been solved

Details

All times are UT, files are luci2.20140117.iiii

Darks for Kurk, 10 * 300sec, #3-12

00:57 Sunset, some cirrus around, not flats, low wind, hum = 15%, T = 2deg

01:00 We are delayed a bit since Jenny is doing some tests for LUCI1

01:45 Steve checks pointing & colimation, looks OK, seeing around 0.6"

Center of rotation

02:00 Pointing to a astrometric star in StoneC onaxis, check pointing & collimate, images on-axis

Tests LS-acq. 2stars with small rotation

02:50 BS9107_longslitACQ_PA0_900_bino, preset in bino-mode, modified script by Alexander,
Steve authorizes LUCI1 before, didn't work since script had a preset and offset as 1st obsitem
remove authorization for LUCI1

02:58 BS9107_longslitACQ_PA0_900, mono-mode, doesn't work

03:00 Alexander restartet tel.service, than preset again, src #26, LS #27, aligned
verification #28/29, nada, slitpos not indicated, redo, src #30, LS #31, aligned verfication #32/33
gna, script moved star back to orig. PA, once more src #34, LS #35, alignment & verification, #36/37, star1 at rot of 7deg OK, but translational offset not, stars are off by dx=-19, dy=-7pix of requested pos.

03:45 TNO 1998WW31, J #38-42,

04:01 Stopped/restarted telescope, restart rotator (Steve)

04:03 Restartet TNO 1998WW31, J #43-51 seeing degraded (0.8" on images, 1-1.4" on guider)

*04:20 #50, error message Aladin "failed to load file, renaming of XFITS-file failed, expected"*
fits instead of xfits-stream. Please fix header

04:25 Ks #52-60, seeing ~0.8" (1-1.5" on guider)

04:40 Stopped/restarted telescope

04:49 BS9107, point and collimate, do field alignment on target (off by dx,dy=-7,-4pix) and try LS-acq w 2stars again
src #67, LS #68, aligned, verification #69/70, pos star1: 891/1135, requested: 1021/1134,
Offset applied: dx,dy=13.6,-1.3",drot=7.7deg, found at: 1007/1125, off by dx,dy=-13,-9pix, seeing 1.2"
do a second round: #71 OK

05:23 Ima-Mov, M67, AM=1.3, PA=0, #72-77 PA=90, #78, bino-preset , worked, readout simultaneously worked too,
PA=180/20, #79/80 all bino-preset,

05:57 Telescope service hung at PA=180 preset, Tom is inspecting the cause,
he seems to have found the cause and fixed it...lets see

Ima stresstest

06:08 Steve's pointing and collimating close to Orion, got an RPC-communication error

06:16 HH91_bino, Michelle put a huge effort in preparing this script, do only one obsitem to see if GS is available, #81 works!!!

06:21 HH91_bino, full set as a benchmark, start with FeII, 36 obsitems (24 on source, 12 on sky), 9*10 integrated, seeing is not good so scientific use extremely limited, but good for stability check, #82-85 had to stop after 4th exp. (#85)
Steve fixed secoondary (run into limits), restart, FeII #86-89, doesn't work, problem is large offsets in bino-mode

06:52 Steve authorized LUCI2 only, restart HH91 (w/o bino), FeII #90-101 , seeing on images ~0.7"

07:25 Another telescope service stop - updating software version potentially will solve problem, Tom and Alexander working on implementation.

07:36 Starting the script again. restart HH91 (w/o bino) at place where we left off, FeII #102-125 , seeing on images ~0.7",
cont'd with H2 (looks great), low AM now but seeing ~0.5" continue until the bitter end (stresstest), #126-139
ended with 3 sky-frames, eve at AM=2.3 still 0.6" seeing

08:58 Ima-Mov, M67, AM=1.3, PA=0,90,180,270, #141-144, Steve check pointing and collimation before

09:08 OB execution panel does not accept commit, restart, OK, but than tel.service hangs (the latter a consequence of the former, ie not a separate issue)

Preparation active flexure compensation

09:34 BS9142, N3.75 slitview ACTflex #145, N1.8 slitview ACTflex #146, src #147, aligned, verification #148, seing 0.5" on guider
N3.75, 210K@2.2mue, 50*3sec w/o flexure compensation, #149-198
N3.75, 200HKK@1.17mue, 50*3sec w/o flexure compensation, #199-248
N1.8, 200HKK@1.17mue, 50*3sec w/o flexure compensation, #249-298
N1.8, 210K@2.2mue, 50*3sec w/o flexure compensation, #299-348

Deep spectroscopy

11:40 Kurk GDNPHB31, sky #349, src #350, MOS-view #351, verification #352, slight correction to account for distortion #353/354,
error FW2, fixed, checked alignment again #355/356, spectra #356-378, last spectrum taken at 13:30, seeing on guider 0.6-0.7"

12:50 - another telescope service problem. stopping and restarting telescope service at #371. continuing script.

13:36 Telluric for GDNPHB31, acq.+ spectra 379-396, some of the last spectra are not useful,
did a good set of Telluric at the requested slits

14:10 Closing up

Calibrations for GDNPHB31

Mask GDNPHB31, G210@2.225mue, clear+K, CalUnit in, all images LIR, all normal
Lamp off, 5*2.75s, #397-401
Halo2 on, 5*20sec, #402-406
Lamps off, 5*20sec, #407-411
Argon on, 1*10sec, #412
Neon on, 1*10sec, #413
Xenon on, 1*10sec, #414
Lamps off, 5*10sec, #415-419

-- JochenHeidt - 19 Jan 2014
Topic revision: r12 - 23 Jan 2014, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback